Kyky232:

Triangle A″B″C″ is formed using the translation (x + 0, y + 2) and the dilation by a scale factor of 2 from the origin. Which equation explains the relationship between segment AC and segment A double prime C double prime? (pic in replies) segment AC over segment A double prime C double prime = 2 segment A double prime C double prime over segment AC = one half segment AC = segment A double prime C double prime over 2 segment A double prime C double prime = segment AC over 2

Kyky232:

1 attachment
jimthompson5910:

AC is 6 units long as the diagram indicates The scale factor 2 means when we dilate, the length of A''C'' will be 2*6 = 12 units long So, A''C'' = 2*AC

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 AC is 6 units long as the diagram indicates The scale factor 2 means when we dilate, the length of A''C'' will be 2*6 = 12 units long So, A''C'' = 2*AC \(\color{#0cbb34}{\text{End of Quote}}\) so D?

jimthompson5910:

We can divide both sides by 2 to get \(A''C'' = 2*AC\) \(\frac{A''C''}{2} = \frac{2*AC}{2}\) \(\frac{A''C''}{2} =AC\) \(AC = \frac{A''C''}{2}\) This means that if we know how long A''C'' is, then we divide by 2 to get the length of AC

Kyky232:

IS IT D OR NOT-

Kyky232:

im just getting more confused-

jimthompson5910:

no, the answer isn't D

Kyky232:

C?

jimthompson5910:

yes